OpenStudy (anonymous):

solve and check. 3d^2+8d-3=0 d=

OpenStudy (anonymous):

3d2+9d-d-3=0 3d(d+3)-(d+3)=0 (3d-1)(d+3)

OpenStudy (anonymous):

d= 1/3 or -3

OpenStudy (anonymous):

now subsititute those two values in equation above........

OpenStudy (anonymous):

I dont know how to do it

OpenStudy (anonymous):

so is the answer 1/3?